### v2

# Prosody Analysis of Audiobooks

Abstract
Recent advances in text-to-speech have made it possible to generate natural-sounding audio from text. However, audiobook narrations involve dramatic vocalizations and intonations by the reader, with greater reliance on emotions, dialogues, and descriptions in the narrative. Using our dataset of 93 aligned book-audiobook pairs, we present improved models for prosody prediction properties (pitch, volume, and rate of speech) from narrative text using language modeling. Our predicted prosody attributes correlate much better with human audiobook readings than results from a state-of-the-art commercial TTS system: our predicted pitch shows a higher correlation with human reading for 22 out of the 24 books, while our predicted volume attribute proves more similar to human reading for 23 out of the 24 books. Finally, we present a human evaluation study to quantify the extent that people prefer prosody-enhanced audiobook readings over commercial text-to-speech systems.
1 Introduction
Audio books are actor-narrated recordings of written texts, typically full length novels. Audio books have grown rapidly in popularity in recent years, with annual sales in the United States reaching \$1.3 billion dollars in 2020.111https://www.statista.com/topics/3296/audiobooks/. The National Library Service for the Blind has recorded tens of thousands of talking books for the sight impaired since 1933, read by an extensive network of volunteers.
Text-to-speech (TTS) technology has made it possible to create machine-generated readings of text. With recent advances in the use of deep learning for TTS models (Oord et al., 2016; Arık et al., 2017), the ability to produce natural-sounding audio closer to human speech has also improved. But there remains a sizable performance gap, especially in the domains like story narration which depend on a global interpretation of a book-length text. Audio books are often recorded by professional actors, involving dramatic vocalizations and intonations by the reader to emphasize emotions, dialogues, and environment and character descriptions in the narrative. Although machine-generated audio sounds human-like in nature, it does not account for these elements in narrative readings, which show much variation in intonation and prosody.
In this paper, we analyze audio books from an NLP-perspective, with two distinct objectives in mind. First, we study how higher-order NLP analysis of novels (e.g. character identification, quote/dialog analysis, and narrative flow analysis) might be used to build better TTS systems.
Second, we are interested as audio books as a source of annotations for narrative texts. One of the challenges of NLP on book-length documents is the cost of annotation: reading a novel is a 10-15 hour commitment, making it cost prohibitive to do human annotation for special-purpose tasks on a large corpora of books. We see human-recorded audio books as a potential solution here: each human recording implicitly contains information about the contents of the texts, which analysis of the audio may enable us to programmatically extract.
State-of-the-art TTS systems do not approach the dynamic range and expressiveness of human readers. Machine-generated audio consistently stays within a small range of pitch and volume, indicative of a flatter reading. In this work, we address the task of generating higher-quality audiobooks closer to human readings in nature, by predicting prosody attributes from the text. Our main contributions in this work include:
  • Character-level analysis of audiobook reader behaviors — We quantify the extent to which the gender properties of characters are reflected in audiobook readings. Specifically, in 21 of 31 books where the two lead characters differ in gender, readers used lower pitch and higher volume to portray the male character. In 52 of 62 books with significant dialog for both genders, reader pitch was higher for female characters. Further, readers generally use lower pitch in narrative regions rather than dialog, independent of gender.
  • Models for audiobook prosody prediction — We address the task of predicting prosody attributes given a narrative text, training a variety of models on audiobook readings to predict prosody attributes (pitch, volume, and rate) at the sentence-level and phrase-level. Best results for all three prosody attributes came using a LSTM architecture, built upon MPNet embeddings, operating at the phrase level. Our predicted prosody attributes correlate much better with human audiobook readings than results from a state-of-the-art commercial TTS system (Google Cloud Text-to-Speech). The dot plot of Figure 1 presents the human-TTS correlation for two systems, on a test set of the first chapters of 24 books. Our predicted pitch attribute shows a higher correlation with human reading for 22 out of the 24 books, while our predicted volume attribute proves more similar to human reading for 23 out of the 24 books.
    (a) Pitch
    (b) Volume
    Figure 1: Enhancing TTS with our predicted prosody attributes show better correlations with human reading than unaided Google Cloud Text-to-Speech.
  • Human evaluation study for text-to-speech audiobook reading — The prosody attributes predicted by our models from text can be directly incorporated to enhance TTS systems through the Speech Synthesis Markup Language (SSML). We conduct a human evaluation study to understand whether humans prefer to hear our more expressive SSML-enhanced text-to-speech audio over the commercial Google Cloud Text-to-Speech, a strong baseline. Results here are inconclusive, but a small majority (12 of 22) subjects preferred our SSML-enhanced readings.
Over the course of this study we have developed a dataset of phrase-aligned books and audiobooks, comprising 1806 chapters across 93 novels, with their corresponding human-read audiobooks and phoneme-level alignment between the text and audio. This dataset will be made publicly available at www.anonymous.com upon acceptance.
This paper is organized as follows. Section 2 presents related work associated with text-to-speech synthesis, including an overview of the Speech Synthesis Markup Language (SSML) we use in our experiments. Section 3 describes our methods for the critical phase of alignment between the textual representation of novels and associated audio book recordings. Section 4 presents our analysis of reader behavior, quantifying the degree to which voices change with a character’s gender. We develop prosody prediction models from text in Section 5, culminating in a human preference study reported in Section 6. Future directions for our work are discussed in Section 7.
2 Related Work
NLP for books:
Recent work focused on performing NLP tasks on novels (Bamman et al., 2014; Brooke et al., 2015; StonyBook, n.d.) has facilitated the analysis and downstream use of book artifacts (Pethe et al., 2020; Kim et al., 2020; Wilmot and Keller, 2020, 2021; Papalampidi et al., 2020; Otake et al., 2020). These systems employ CoreNLP (Manning et al., 2014), Stanza (Qi et al., 2020), and SpaCy (Honnibal and Montani, 2017) for NLP annotation tasks including tokenization, lemmatization, part-of-speech tagging, dependency tagging, named entity recognition, coreference resolution, and quote extraction and attribution.
Approaches to character identification in novels involves clustering variants of character names (Davis et al., 2003; Elson et al., 2010; Vala et al., 2015; Jahan and Finlayson, 2019). (Groza and Corde, 2015) use a folktale-specific ontology based approach. The CHARLES tool (Vala et al., 2016) provides a collaborative interface for the manual annotation of characters. (Dekker et al., 2019) compare four different approaches for character identification by evaluating against a manually annotated dataset. Further, methods to generate character embeddings and analyze relationships between characters (Inoue et al., 2022; Iyyer et al., 2016; Xanthos et al., 2016; Azab et al., 2019; Brahman et al., 2021) provide additional information about characters apart from their mentions in the book.
Genre identification (Worsham and Kalita, 2018), tracking emotion in narratives (Mohammad, 2011), and quote identification and attribution (O’Keefe et al., 2012; Zhang et al., 2003) are additional tasks that provide context about events occurring in the book narrative.
These artifacts obtained from analysis of books contain information that influences the way the readings of these books can be customized as per the type of content, mood, characters’ nature, and context of dialogue.
Text to Speech:
Speech synthesis or text-to-speech (TTS) is the technology to produce artificial human speech. With recent advances in deep learning, neural network-based TTS models have become the dominant approach. WaveNet (Oord et al., 2016) is a neural vocoder that directly generates time domain waveforms from linguistic features without the use of acoustic features, and it outperforms traditional statistical parametric synthesizers. Tacotron(Wang et al., 2017) uses a single neural network to directly generate spectrograms from a sequence of characters without the use of linguistic features. (Shen et al., 2018) extend this work by combining it with a modified WaveNet to improve voice quality. Naturalness and robustness of speech synthesis are improved based on this work by transferring prosody and speaking style from reference signals (Skerry-Ryan et al., 2018), using reference encoder and style tokens (Wang et al., 2018), hierarchical latent variables (Hsu et al., 2019), multi-reference encoder (Bian et al., 2019), and duration predictors (Yu et al., 2019; Shen et al., 2020). (Weiss et al., 2021) propose Wave-tacotron, an end-to-end model that generates waveform directly from text. Tacotron model is further improved by changing sequence generation from autoregressive to non-autoregressive (Elias et al., 2021b, a).
DeepVoice (Arık et al., 2017) is a CNN-based model that uses independent neural networks in place of traditional TTS components. (Arik et al., 2017) extend this work by adding multi-speaker support using Tacotron-style model. DeepVoice 3 proposed by (Ping et al., 2017) has a fully-convolutional attention-based encoder-decoder architecture that supports larger datasets and faster training. FastSpeech (Ren et al., 2019) is a non-autoregressive Transformer-based parallel TTS. It uses knowledge distillation and is able to fully utilize GPU and synthesize spectrograms faster than autoregressive models. Its monotonic constraint on alignment also reduces synthesis failure. FastSpeech 2 is proposed by (Ren et al., 2020), which solves the one-to-many mapping problem better by conditioning on supervised fundamental frequency and energy. FastPitch (Łańcucki, 2021) uses a similar approach. Lightweight TTS models (Luo et al., 2021; Nguyen et al., 2021; Vainer and Dušek, 2020) are efficient to deploy on mobile and edge devices.
More recent works on TTS include Pitchtron (Jung and Kim, 2020), which proposes two prosody transfer models to generate speech with rich prosody. (Sini, 2020) investigates different characteristics that contribute to the expressiveness of a voice. (Sivaprasad et al., 2021) present a TTS system that allows user to generate speech with chosen emotion from a continuous and interpretable space of arousal and valence.
Some automated audiobook generation vendors include DeepZen, Speechki, Speechkit, Descript, Pozotron, Scribe Audio, Fabula, Respeecher, and Accessibility.
2.1 SSML Overview
Speech Synthesis Markup Language (SSML) (Taylor and Isard, 1997) is a markup language that provides a standard way to mark up text for the generation of synthetic speech, allowing for more fine-grained control over the generated audio. In this work, we are specifically interested in the <prosody> tag, which can be used to control three different attributes: pitch, volume, and rate of speech:
Figure 2: Sample SSML text with custom prosody attributes.
  • Pitch — The pitch can be increased or decreased by N percent relative to the default value, by setting the pitch attribute as "+N%" or "-N%" respectively.
  • Volume — The volume can be increased or decreased by N decibels relative to the default value, by setting the volume attribute as "+NdB" or "-NdB" respectively.
  • Rate — The rate of speech can be specified as N percent of the default value by setting the rate attribute to "N%".
3 Audio/Text Alignment Methods
With the end goal of enhancing prosody in automatically generated audiobooks, we present a dataset of pairs of books and their corresponding audiobooks from the Project Gutenberg dataset, along with the extracted prosody data from the human-read audiobooks.
The dataset contains a total of 1806 chapters across 93 books, with a mean of 19.42 chapters per book. Generally each Gutenberg audiobook is read by a single narrator, but there are exceptions, so we performed our analysis on a per-chapter basis. We used a mix of narrator annotation and audio speaker identification techniques to establish narrator consistency.
We perform the following data processing steps to generate aligned data and extract the audio attributes.
  • Transcription — We first generate a text transcript of each audiobook. We split the audiobook into fragments using silence for segmentation. For each segment, we attempt to generate a transcript using Google Speech Recognition  [^2] https://pypi.org/project/SpeechRecognition/. Note that this transcript is noisy, and cannot be used directly for further tasks. We use this noisy transcript to align with the appropriate chapter, and use this text for downstream tasks.
  • Alignment — We use the Gentle forced aligner [^3] https://github.com/lowerquality/gentle to generate alignments between text and audio, including timestamp and phoneme information at word-level granularity.
  • Audio Attribute Extraction — To extract the pitch (fundamental frequency in Hz) and the volume (intensity in dB), we use the Python library Parselmouth [^4] https://parselmouth.readthedocs.io/en/stable/, with the timestep set to 0.01 second (compatible with the granularity of the alignment output from Gentle).
    To extract the rate of speech, we perform the following steps: For each phoneme, we first compute the z-scores of the durations of all occurrences. Then for each audio segment (phrase / sentence), we compute the mean of the phoneme duration z-scores, and compute the z-score of the resultant means.
  • Sentence Segmentation — For each sentence, we generate a constituency parse tree using the Berkeley Neural Parser (Kitaev and Klein, 2018) with SpaCy (Honnibal and Montani, 2017). We then split the sentence into phrases, using the constituent sentence (S) components and punctuation.
4 Audiobook Reader Behavioral Analysis
(a) Pitch
(b) Volume
Figure 3: Pitch and volume breakdown for the primary and secondary characters in each book, colored by gender pairs.
We also compare pitch and volume for the most frequent and the second most frequent characters in each book. Figure 3(b) shows pitch and volume of character pairs colored by gender. When the two main characters are of opposite gender, female character has a higher pitch and a lower volume compared to male character as most red dots are below the regression line, and most purple dots are above on both plots. As shown in Table 1, when the two main characters are of opposite gender, we have a 67.74% chance of distinguishing female from male based on pitch or volume.
Prosody
Gender of 1st and
2nd character
1st > 2nd 1st < 2nd
Pitch Female, male 13 5
Male, female 5 8
Male, male 19 18
Female, female 8 5
Volume Female, male 6 12
Male, female 9 4
Male, male 22 15
Female, female 4 9
Rate Female, male 8 10
Male, female 5 8
Male, male 17 20
Female, female 9 4
Table 1: Number of character pairs above and below the regression line for pitch, volume, and rate of speech.
Finally, we compare the pitch and volume for dialogue utterances averaged separately across all the male and female characters in a book. For this comparison, we restrict our analysis to 62 books in which male and female characters each have at least 100 combined word utterances.
Figure 4: Mean pitch for dialog of male and female characters. Female characters are voiced with higher pitch.
For 52 out of 62 books, the mean pitch is higher for female character dialogue as compared to male (binomial p=3e-8). For 32 out of 62 books, the mean intensity (volume) is higher for male character dialogue as compared to female (binomial p=0.449).
5 Predicting Prosody Attributes
We now address the task of predicting prosody attributes (pitch, volume, and rate) for each sentence or phrase, given a chapter text as input. We use Mean Squared Error (MSE) between human readers’ actual prosody values and our predicted prosody values as accuracy measurement.
5.1 Text Input Representations
We experiment with three input embeddings:
  • TF-IDF representations: We use TF-IDF vectorization, limiting the vocabulary to 1,000 words to generate 1000-dimensional representations of the input text.
  • Mean-pooled GloVe embeddings: We compute the 300-dim average word embedding of the input text using the CommonCrawl pre-trained GloVe model (840B tokens).
  • MPNet embeddings: We compute 768-dimensional embeddings of the input text using MPNet (Song et al., 2020) (all-mpnet-base-v2).
We use Sentence-Transformers  [^5] https://www.sbert.net/ to compute the GloVe and MPNet embeddings. We process the audio attributes extracted as described in Section 3 for this task. For each text segment in the input, we consider the values for the three prosody attributes (pitch, volume, rate) in the corresponding audio segment. We compute a single value for each attribute per text segment, and compute the z-scores across all segments in the chapter.
5.2 Models
We use a train-test split of 75-25% across books, i.e. we use 69 books (1,392 chapters) in the training set and 24 books (414 chapters) in the test set. All the models are trained on a 2GHz CPU.
5.3 Sentence-level Attribute Prediction
We perform the prediction task at the sentence-level granularity. In this section, we present the test metrics for baseline models, where we use only a single sentence representation as the input, and predict a single prosody attribute (pitch/volume/rate) as the output.
Embedding Model Pitch Volume Rate
TF-IDF LinReg 0.9217 0.9054 0.8439
MLP (5, 5) 0.9165 1.0000 1.0000
MLP (10, 10) 0.9636 0.9342 0.8034
MLP (20, 20) 1.0554 1.0127 0.8605
GloVe LinReg 0.9103 0.8900 0.8578
MLP (5, 5) 0.9088 0.8955 0.8212
MLP (10, 10) 0.9208 0.8967 0.8396
MLP (20, 20) 0.9749 0.9395 0.8734
MPNet LinReg 0.8667 0.8440 0.7811
MLP (5, 5) 0.8785 0.8457 0.7996
MLP (10, 10) 0.9403 0.8888 0.8519
MLP (20, 20) 1.0642 0.9742 0.8899
Table 2: Test MSE for sentence-level prosody attributes
Table 2 shows the test metrics for the respective baseline models. As the sentence embedding quality improves, we expect the error to reduce. The test metrics are in accordance with this, as MPNet embeddings have the least prediction error across all three attributes.
5.4 Phrase-level Attribute Prediction
A drawback of the sentence-level model is the inconsistency in sentence length. Sentences may range from 1 to more than 500 tokens. In order to introduce consistency, we use phrase representations as input. We segment sentences into phrases using the method described in Section 3.
Table 3 shows the test metrics for the phrase-level baseline models. These models also show similar trends in the results, with MPNet embeddings showing the best performance.
As the MSE for the phrase-level models is lower than that of the sentence-level models, we present results from the phrase-level models subsequently.
Embedding Model Pitch Volume Rate
TF-IDF LinReg 0.9217 0.8789 0.8439
MLP (5, 5) 0.9086 0.8594 0.8145
MLP (10, 10) 0.9104 0.8623 0.8058
MLP (20, 20) 0.9155 0.8700 0.8073
GloVe LinReg 0.9197 0.8644 0.9225
MLP (5, 5) 0.9095 0.8462 0.8717
MLP (10, 10) 0.9037 0.8449 0.8597
MLP (20, 20) 0.9123 0.8496 0.8569
MPNet LinReg 0.8803 0.8048 0.8123
MLP (5, 5) 0.8661 0.7812 0.7763
MLP (10, 10) 0.8675 0.7814 0.7791
MLP (20, 20) 0.8955 0.7933 0.7987
Table 3: Test MSE for phrase-level prosody attributes
5.5 Joint Prediction
Here we present the test metrics for various models at the phrase level, where we train a single common model to predict all three prosody attributes.
Using only a single phrase embedding as input (with no contextual information about the neighboring phrases), we predict the three attributes. Table 4 shows the test MSE for the three attributes. These models show a slight improvement over the best model results from Section 5.4.
Embedding Model Pitch Volume Rate
TF-IDF LinReg 0.9217 0.8789 0.8439
MLP (5, 5) 0.9103 0.8661 0.8264
MLP (10, 10) 0.9080 0.8609 0.8106
MLP (20, 20) 0.9101 0.8618 0.8067
GloVe LinReg 0.9197 0.8644 0.9225
MLP (5, 5) 0.9173 0.8604 0.8905
MLP (10, 10) 0.9120 0.8488 0.8695
MLP (20, 20) 0.9014 0.8474 0.8615
MPNet LinReg 0.8803 0.8048 0.8122
MLP (5, 5) 1.0001 1.0000 1.0000
MLP (10, 10) 0.8667 0.7817 0.7733
MLP (20, 20) 0.8668 0.7798 0.7666
Table 4: Test MSE for non-contextual joint prediction of prosody attributes
5.5.1 Contextual Prediction
We augment the feature space with embeddings of context phrases before and after each phrase, and jointly predict the three attributes.
Embedding Model Pitch Volume Rate
GloVe LinReg 0.9121 0.8513 0.9154
MLP (5, 5) 0.9081 0.8422 0.8836
MLP (10, 10) 0.9056 0.8387 0.8645
MLP (20, 20) 0.9070 0.8439 0.8856
MPNet LinReg 0.8620 0.7779 0.7813
MLP (5, 5) 0.8659 0.7772 0.7852
MLP (10, 10) 0.8594 0.7671 0.7751
MLP (20, 20) 0.9381 0.7831 0.8586
Table 5: Test MSE for contextual joint prediction of prosody attributes
Table 5 shows the test MSE for the three attributes. This is again a slight improvement over non-contextual prediction (Table 4).
5.5.2 Sequential Prediction
We use a neural network (LSTM architecture and training described in Section 5.2) to predict the prosody attributes for all input. We experiment with sequence lengths 2 and 3.
Embedding Model Pitch Volume Rate
GloVe LSTM(len 2) 0.8897 0.8260 0.8358
MPNet LSTM(len 2) 0.8387 0.7540 0.7488
LSTM(len 3) 0.8362 0.7518 0.7449
Table 6: Test MSE for joint prediction of prosody attributes with LSTM
Table 6 shows the test MSE for the three attributes. The LSTM model trained on MPNet phrase embeddings, with sequence length 3 shows the best performance. We use this model further downstream for text-to-speech audio generation and human evaluation.
5.5.3 Prediction on Character Dialogue
We observe that human readers read character dialogues more expressively than descriptive text.
Figure 5: Mean ground-truth pitch z-score for quote and non-quote phrases from chapters in the test set
Figure 5 shows the distribution of the mean pitch z-scores for phrases that contain quotes or are a part of quotes, and phrases that do not contain quotes. Readers tend to use a higher pitch when reading dialogue and a lower pitch when reading descriptive text, as shown by the right and left shifts from 0 in the distribution.
Figure 6: Mean predicted pitch z-score for quote and non-quote phrases from chapters in the test set
Figure 6 shows the same distributions for our model predictions. These follow a similar pattern, indicating that the model captures information about dialogue, even though we do not explicitly use this information during training.
We further examine this effect by looking at the prosody attribute prediction on dialogues only. Table 7 shows the MSE on a subset of the test data containing only dialogues. Pitch and volume attributes on dialogues are predicted with lower error compared to all input.
Embedding Model Pitch Volume Rate
GloVe LSTM(len 2) 0.8568 0.7683 0.7773
MPNet LSTM(len 2) 0.8073 0.6881 0.8291
LSTM(len 3) 0.8018 0.6872 0.8271
Table 7: Test MSE for joint prediction of prosody attributes on dialogues
Character Embeddings:
More subtle information about the nature of individual characters (e.g. hero vs. villain, old vs. young) inform actor’s readings. We conducted an experiment to augment our models by concatenating 908d character embeddings (Inoue et al., 2022) to text representations before applying PCA for dimension reduction.
Embedding (Size) Model Pitch Volume Rate
Baseline N/A 1.3543 1.9550 2.3218
GloVe + CE (200) LSTM(len 3) 1.0573 0.9559 1.2786
GloVe + CE (100) LSTM(len 3) 0.9181 0.8021 1.3043
GloVe + CE (50) LSTM(len 3) 0.9163 0.8169 1.4364
MPNet + CE (200) LSTM(len 3) 0.9384 0.8368 1.2711
MPNet + CE (100) LSTM(len 3) 0.8736 0.7631 1.0610
MPNet + CE (50) LSTM(len 3) 0.8734 0.7709 1.0860
Table 8: Test MSE for joint prediction of prosody attributes on dialogues using character embeddings (CE) different PCA dimensions.
Table 8 shows the MSE on plain audio generated without SSML attributes and enhanced audio generated with SSML attributes predicted from text and character embeddings. The results improve over our baseline, but not not over our best models.
6 Generating and Evaluating Text-to-Speech Audiobooks
We use the best performing model (LSTM with sequence length 3, and MPNet embeddings as input) to generate predictions for each book in the test set. We use a sliding window for inference, and compute the final prediction for each text input as the mean of its prediction in all windows.
To convert the predicted z-scores into SSML attributes, we compute the mean and standard deviation for the pitch and volume in each human-read audio file in the test set, and use these values to convert the prosody attributes into absolute values (Hz for pitch and dB for volume). Using the mean pitch/volume as reference, we convert these into relative values as required by the SSML specification (as described in Section 2.1). For rate of speech, we use a fixed value for mean and standard deviation (100 and 50 respectively) and perform the same computation. Note that we use these values as we have a reference point from the human-read audios, but these can be adapted as per the TTS voice to introduce more prosodic variations.
6.1 Human Evaluation
We have demonstrated that our pitch, volume, and rate models successfully predict the prosody of human audiobook readers. However, it is not apparent whether human listeners will prefer text-to-speech synthesis augmented by our analysis over SSML models developed by commercial TTS vendors. This is a very difficult baseline for us to try to beat, given the vast training resources deployed to build commercial TTS systems.
We performed a Mechanical Turk study to evaluate our models. For each of the 24 books in the test set, we identify the phrases with the highest absolute values of predictions across the three prosody attributes. We then generate audio for this phrase text along with one context phrase before and after, using the Google Cloud Text-to-Speech API  [^8] https://cloud.google.com/text-to-speech with the voice en-US-Standard-A. We produce two different audio variants for each text:
  • Plain: We enclose the text in a <speak> tag, without additional prosody information.
  • SSML: We enclose each phrase in a <prosody> tag with the three predicted attributes specified: see example in Figure 2.
For each pair of audios, we ask five annotators on Amazon Mechanical Turk to rate which audio sounds better and more natural with intonations similar to human readers.
SSML Plain Neutral Count
5 0 0 1
4 1 0 2
3 1 1 2
3 2 0 7
2 2 1 2
2 3 0 5
1 3 1 1
1 4 0 2
0 5 0 2
Table 9: Annotator preference counts: denotes whether majority annotators preferred SSML or Plain text.
Table 9 shows the results of the evaluation. Each row shows the annotator preferences, and the number of books from the test set with this distribution. For example, row 2 means that there are 2 pairs of audios for which 4 annotators preferred the SSML-enhanced version over the plain text version. For 12 out of 24 test sample pairs, a majority of the annotators preferred the SSML-enhanced version.
Qualitative Analysis:
7 out of the 10 pairs in which the majority preferred plain text had phrases with higher values for pitch and volume in the SSML-enhanced version. Also, for 8 of 12 pairs in which the majority preferred the SSML version, the rate of speech was lower than the default.
7 Conclusion
We have developed an aligned text-audio book corpus, and deployed it to study both reader behavior and improve prosody models for TTS generation of audiobooks. We have shown that actors customize voices to represent specific characters, and adopt more expressive voices than those generated by conventional TTS systems, motivating the development of full-book analysis methods to understand/predict a reader’s goals in narrating a text.
Limitations
  • Our approach produces expressive audio through a sequence-to-sequence model producing annotated SSML to feed to commercial TTS system, instead of directly-trained neural text-to-speech models (text in, audio out). Presumably higher performance scores would follow from the more direct approach, given enough training data.
  • Our sequence-to-sequence approach makes it difficult to control expressiveness at a low level of granularity: word-level prosody cannot be properly controlled because the generated speech sounds disconnected. We believe this is why our models did not perform better at human evaluation.
  • Our research relies on paid access to a commercial TTS system, which limits the size of our experiments. Available open source TTS systems produced significantly poorer speech, make it less attractive as a foundation for our work.